OpenStudy (anonymous):

What is the x-coordinate of the solution of the following system of equations? 3x + y = 6 x - y = 6

OpenStudy (anonymous):

3x + y = 6 = y =-3x+6 x - y = 6 x-(-3x+6)=6 x+3x-6=6 4x+6=6 -6 -6 4x=0 x=0

OpenStudy (anonymous):

I think i'm wrong. Let me recheck. :S

OpenStudy (anonymous):

1st one is wrong.

OpenStudy (anonymous):

Im going to go on a limb and say that the 0 answer is correct

OpenStudy (anonymous):

No it cant be the answers to my course say -3 0 3

OpenStudy (anonymous):

3x + y = 6 = y =-3x+6 x - y = 6 x-(-3x+6)=6 x+3x-6=6 4x-6=6 +6 +6 4x=12 x=3

OpenStudy (anonymous):

I see where i made the mistake.. In the 1st one that +6 was suppose to be -6

OpenStudy (anonymous):

u see?

OpenStudy (anonymous):

Ahhh

OpenStudy (anonymous):

yea i do

OpenStudy (anonymous):

ahah, so 3.. cya

OpenStudy (anonymous):

thanks